$2500 is to be deposited in Banks A and B. The interest rate per annum for Bank A is 6% while that for Bank B is 6,5%. After one year, their interests are equal. How much money is to be deposited in each bank?
Answer by richwmiller(17219)

x+y=2500
6/100*x=6.5/100*y
0.06*x=0.065*y
x+y=2500
We substitute for x
0.06*(2500-y)=0.065*y
We multiply out
0.06*2500-0.06y=0.065*y
We combine like terms.
1500=1.25*y
Isolate y
y=1200.0 at 6.5%
Calculate x
x+y=2500
x=1300 at 6%
Now,we know the total invested is: 2500
Total invested $1300+$1200=$2500
We check
0.06*1300=0.065*1200
78=78.0
ok